OpenStudy (anonymous):

Bob has 5 more dimes than quarters. He has $3.30. How many of each coins does he have.

OpenStudy (saifoo.khan):

i did one before!

OpenStudy (anonymous):

On my other question? Its not appearing.

OpenStudy (saifoo.khan):

He got 13 dimes, and 8 quatrers!

OpenStudy (anonymous):

And how did you get to that?

OpenStudy (anonymous):

He has 13 dimes, and 8 quarters.

OpenStudy (anonymous):

And someone else already answered....

OpenStudy (saifoo.khan):

Lol, and someone copied! jk jk

OpenStudy (anonymous):

Haha but how did you get to your answer ?

OpenStudy (saifoo.khan):

\[\Huge My-Mind-works!\]

OpenStudy (saifoo.khan):

\[means..... \Huge Mentally!\]

OpenStudy (anonymous):

this doesnt help much. lol

OpenStudy (anonymous):

I got to my answer by just being intuitive. Meaning I just thought about it and even counted on my fingers lol. Then I got my answer.

OpenStudy (anonymous):

It can be solved with algebra. Number of quarters = q and number of dimes = d. We have, from the question: d = q+5 (there are five more dimes than quarters), and 10d + 25q = 330 (total amount of money = $3.30, or 330 cents) Putting them together: 10(q+5) +25q = 330 35q + 50 = 330 35q = 280 Therefore q=8, and d=13, as stated above.